OpenStudy (eric_d):

Given that tan theta=0.5744, find all the values of theta in range of...................., giving ur answers correct to nearest 0.1 degree

OpenStudy (eric_d):

\[0 \le \theta \le 360\]

OpenStudy (eric_d):

@Hero

OpenStudy (eric_d):

Nvm, I managed to solve it